# Leveraging Offline Data in Linear Latent Bandits

### v1

@@ 5 Using Offline Data to Speed Up Online Learning with LOCAL-UCB
1:  Input: Projection matrix $\hat{\textbf{U}}\hat{\textbf{U}}^{\top}$ and confidence bound $\Delta_{\operatorname{off}}$ from an offline uncertainty-aware method, e.g. SOLD.
@@
$\textstyle{a_{t},\widetilde{\bm{\beta}}_{t},\widetilde{\textbf{U}}_{t}\leftarrow\arg\max_{a,\bm{\beta},\textbf{U}}\phi(x_{t},a)^{\top}\bm{\beta}}\text{ such that }$
$\textstyle{\hat{\bm{\beta}}_{1,t}\leftarrow\textbf{U}(\textbf{U}^{T}\textbf{V}_{t}\textbf{U})^{-1}\textbf{U}^{\top}\textbf{b}_{t},\ \|\textbf{U}^{\top}(\bm{\beta}-\hat{\bm{\beta}}_{1,t})\|_{(\textbf{U}^{\top}\textbf{V}_{t}\textbf{U})^{-1}}\leq\alpha_{1,t}}$
@@
$\textstyle{\|\bm{\beta}\|_{2}\leq R,\textbf{U}^{\top}\textbf{U}=\textbf{I}_{d_{K}},\textbf{U}\textbf{U}^{\top}\bm{\beta}=\bm{\beta},\ \|\hat{\textbf{U}}^{\top}\textbf{U}\|_{F}\geq\sqrt{d_{K}-\Delta_{\operatorname{off}}^{2}/2}}$
@@
We provide the following guarantee for LOCAL-UCB. Notice that our guarantee shows that for enough offline data with $N\gg T$, the effective dimension of the problem is $d_{K}$. It increases to $d_{A}$ as $T$ gets closer to $N$. The quality of the offline data is reflected in the coverage constants $\lambda_{D}$ and $\lambda_{\min}$.
@@
Under Assumptions 1 and 2, if $\alpha_{1,t}=R\sqrt{\mu}+CR\sqrt{d_{K}\log(2T/\delta)}$ and $\alpha_{2,t}=R\sqrt{\mu}+CR\sqrt{d_{A}\log(2T/\delta)}$ for a suitable universal constant $C$, with probability at least $1-\delta$, LOCAL-UCB has regret
$\operatorname{Reg}_{T}$ $=O\textstyle{\left(\min\left(Rd_{A}\sqrt{T},Rd_{K}\sqrt{T}\left(1+\frac{1}{\lambda_{\theta}\lambda_{A}^{3}}\sqrt{\frac{Td_{A}}{d_{K}N}}\right)\right)\right)}.$
@@
However, the subspace constraint $\|\hat{\textbf{U}}^{\top}\textbf{U}\|_{F}\geq\sqrt{d_{K}-\Delta_{\operatorname{off}}^{2}/2}$ is nonconvex. In fact, we lower bound a convex function in the constraint. This essentially searches over a complicated star-shaped set in the space of all U. As such, LOCAL-UCB is unfortunately computationally inefficient.
@@
Figure 1: Left: Geometric interpretations of LOCAL-UCB. Showing $\mathcal{C}^{t}(\bm{\beta},\operatorname{on})\cap\mathcal{C}^{t}(\bm{\beta},\operatorname{off})$ in green for three timepoints $t=t_{1},t_{2},t_{3}$. The dotted lines delineate the subspace confidence set. Right: Geometric interpretation of ProBALL-UCB. $\mathcal{C}^{t_{1}}(\bm{\beta},\operatorname{on})\not\subset\widetilde{\mathcal{C}}^{t_{1}}(\bm{\beta},\operatorname{off})$, so we continue to use projections; but by time $t_{2},$ $\mathcal{C}^{t_{2}}(\bm{\beta},\operatorname{on})\subset\widetilde{\mathcal{C}}^{t_{2}}(\bm{\beta},\operatorname{off})$, so we stop using projections.
6 More Practical Optimism with ProBALL-UCB
We note that while LOCAL-UCB is a natural algorithm with promising guarantees, it is not computationally efficient. We address this issue by introducing ProBALL-UCB (Algorithm 3), a practical and computationally efficient algorithm for accelerating online learning in linear latent bandits with offline data. We make three simplifications compared to LOCAL-UCB. Namely, (1) this has cruder confidence sets, (2) we wait for one confidence set to become a subset of the other instead of considering complex intersections, and (3) we use a computable proxy for the above condition instead of explicitly checking it. Surprisingly, this leads to only a minor degradation in provable guarantees.
To geometrically motivate our update rule like in Section 5, we consider the $d_{K}$-dimensional confidence ellipsoid $\mathcal{C}^{t}(\bm{\beta},\hat{\textbf{U}})$ of $\bm{\beta}$ inside subspace $\hat{\textbf{U}}^{\top}$. First, since $\|\hat{\textbf{U}}\hat{\textbf{U}}^{\top}-\textbf{U}_{\star}\textbf{U}_{\star}^{\top}\|_{2}\leq\Delta_{\operatorname{off}}$, we can show that a crude bound for $\phi(x,a)^{\top}\bm{\beta}_{\star}$ is $\max_{\bm{\beta}\in\mathcal{C}^{t}(\bm{\beta},\hat{\textbf{U}})}\phi(x,a)^{\top}\bm{\beta}+\Delta_{\operatorname{off}}$. This corresponds to optimism over a crude confidence ball $\widetilde{\mathcal{C}}^{t}(\beta,\operatorname{off})$ of radius $\Delta_{\operatorname{off}}$ around $\mathcal{C}^{t}(\beta,\hat{\textbf{U}})$. Second, instead of being optimistic over the intersection of this ball with our usual $d_{A}$-dimensional confidence ellipsoid $\mathcal{C}^{t}(\beta,\operatorname{on})$, we merely wait for $\mathcal{C}^{t}(\beta,\operatorname{on})$ to be inside $\widetilde{\mathcal{C}}^{t}(\beta,\operatorname{off})$. Finally, instead of actually checking whether $\mathcal{C}^{t}(\beta,\operatorname{on})$ is inside $\widetilde{\mathcal{C}}^{t}(\beta,\operatorname{off})$, we use a proxy condition that roughly demands $\Delta_{\operatorname{off}}T>d_{A}\sqrt{T}$. Here, $d_{A}\sqrt{T}$ represents the "cumulative size" of the ellipsoid across timesteps $t$, while $\Delta_{\operatorname{off}}T$ represents the "cumulative size" of the ball. This is illustrated in Figure 1.
We note that the idea behind ProBALL-UCB is fairly general and can be used to design adaptations of other algorithms as well. Broadly, we project to a subspace, perform optimism there, and track the cumulative error for using this subspace. Once the cumulative error of using the subspace ($\approx\Delta_{\operatorname{off}}T$ in our case) exceeds the cumulative error of not using the subspace ($\approx d_{A}\sqrt{T}$), we stop using the subspace. Although ProBALL-UCB instantiates this framework with LinUCB, the same idea can be immediately applied to other bonus-based algorithms like SupLinUCB.
@@ 6 More Practical Optimism with ProBALL-UCB
$\operatorname{Reg}_{T}$ $=\textstyle{\widetilde{O}\left(\min\left(Rd_{A}\sqrt{T},Rd_{K}\sqrt{T}\left(1+\frac{1}{\lambda_{A}^{3}\lambda_{\theta}}\left(\sqrt{\frac{d_{A}T}{d_{K}N}}+\sqrt{\frac{d_{A}}{SN}\sum_{t=1}^{S}\kappa_{t}^{2}}\right)\right)\right)\right)}.$
In the worst case, $\kappa_{t}=O(t)$ and so $\frac{1}{S}\sum_{t=1}^{S}\kappa_{t}^{2}=O(T^{2})$, but if all features $\phi(x_{t},a_{t})$ lie in the span of $\hat{\textbf{U}}$ for $t\leq S$, then $\frac{1}{S}\sum_{t=1}^{S}\kappa_{t}^{2}=O(T)$.
@@ 7 Experiments
We now establish the practical efficacy of SOLD (Algorithm 1) and ProBALL-UCB (Algorithm 3) for linear latent contextual bandits through a series of numerical experiments. We perform a simulation study and a demonstration using real-life data. [^4] All experiments were run on a single computer with an Intel i9-13900k CPU, 128GB of RAM, and a NVIDIA RTX 3090 GPU, in no more than an hour in total. We obtain confidence bounds $\Delta_{\operatorname{off}}$ using three different concentration inequalities – (1) Hoeffding as in Proposition 2 (H-ProBALL), (2) empirical Bernstein as in Proposition 1 (E-ProBALL), and (3) confidence intervals generated by the martingale Bernstein concentration inequalities of Waudby-Smith and Ramdas (2023); Wang and Ramdas (2024) (M-ProBALL). We use a simpler expression for $\Delta_{\operatorname{off}}$, set $\tau^{\prime}=0$, and vary the hyperparameter $\tau$ to adjust for overly conservative $\Delta_{\operatorname{off}}$ [^5] Namely, we set $\Delta_{D}=0$ in $\Delta_{\operatorname{off}}$. Also, Lemma 6 and some thought reveal that choosing $\tau^{\prime}=0$ recovers the ”good” version of ProBALL-UCB guarantees, if features are isotropic enough..
Simulation study.
We first perform a simulation study on a latent linear bandit with $d_{A}=50$ and $d_{K}=2$. We generate $\textbf{U}_{*}$ with $\textbf{U}_{ij}\text{ i.i.d. }\text{Unif}(0,\frac{2.5}{d_{K}d_{A}})$. We simulate the hidden labels $\bm{\theta}_{n}\sim\mathcal{N}(0,d_{K}^{-1}\textbf{I}_{d_{K}})$, generate feature vectors $\phi(x_{n,h},a_{n,h})\sim\mathcal{N}(0,\textbf{I}_{d_{A}})$ normalized to unit norm, and sample noise $\epsilon_{n,h}\text{ i.i.d. }\mathcal{N}(0,0.5^{2})$. We use SOLD to estimate $\hat{\textbf{U}}$ from the offline dataset $\mathcal{D}_{\operatorname{off}}$, which consists of $5000$ trajectories of length $20$ each. In accordance with the confidence set determined by Li et al. (2010), we choose $\alpha_{1,t}=0.33\sqrt{d_{K}\log(1+10T/d_{K})}$ and $\alpha_{2,t}=0.33\sqrt{d_{A}\log(1+10T/d_{A})}$, and share the LinUCB and ProBALL-UCB hyperparameters by assigning $\alpha_{t}=\alpha_{2,t}$.
Figure 2: Comparison of ProBALL-UCB with LinUCB, for different choices of $\tau$ and confidence bound constructions. All variants perform no worse than LinUCB, with martingale Bernstein performing the best. The shaded area depicts $1$-standard error confidence intervals over $30$ trials.
@@
The results are presented in Figure 2. Note that ProBALL-UCB (Algorithm 3) performs no worse than LinUCB, no matter what we choose for $\tau$ and $\Delta_{\operatorname{off}}$. However, we see a distinct benefit from using tighter confidence bounds – as $\Delta_{\operatorname{off}}$ gets smaller, Algorithm 3 chooses to utilize the projected estimate $\hat{\bm{\beta}}_{1,t}$ more often, resulting in better performance. Note that the kinks in the regret curves correspond to points where Algorithm 3 switches over to the higher dimensional optimism in Line 9.
MovieLens dataset.
In line with Hong et al. (2020), we assess the performance of our algorithms on real data using the MovieLens dataset (Harper and Konstan, 2015). This is a large-scale movie recommendation dataset comprising $6040$ users and $3883$ movies, where each user may rate one or more movies. Like Hong et al. (2020), we filter the dataset to include only movies rated by at least $200$ users and vice-versa. We factor the sparse rating matrix into user parameters $\bm{\beta}$ and movie features $\Phi$ using the probabilistic matrix factorization algorithm of Mnih and Salakhutdinov (2007a), using nuclear norm regularization so that the rank of $\bm{\beta}$ is $d_{K}=18$. However, we consider a much higher dimensional problem than Hong et al. (2020) do – we let $d_{A}=200$ so $\bm{\beta}\in\mathbb{R}^{1589\times 200},\Phi\in\mathbb{R}^{200\times 1426}$. At each round for user $i$, the agent chooses between $20$ movies of different genres with features $\Phi_{a_{1}},...,\Phi_{a_{20}}$, and has to recommend the best movie presented to it to maximize the user’s rating of the movie. We generate rewards for recommending movie $j$ to user $i$ by $\bm{\beta}_{i}^{T}\Phi_{j}+\epsilon_{ij},\epsilon_{ij}\text{ i.i.d. }\mathcal{N}(0,0.5)$.
Our hyperparameters are chosen and varied just as in the simulation study. To reproduce the methods of Hong et al. (2020), we cluster the user features into $d_{K}$ clusters using k-means, and provide mUCB and mmUCB with the mean vectors of each cluster as latent models. We initialize ProBALL-UCB with a subspace estimated with an unregularized variant of SOLD, that uses pseudo-inverses instead of inverses, because of difficulties in finding an appropriate regularization parameter for this large, noisy, and high-dimensional dataset. The subspace was estimated from $5000$ trajectories of length $50$ simulated from the reward model and the uniform behavior policy. Note that we assign $\Delta_{\operatorname{off}}$ for $\epsilon$ in mmUCB, as this is their tolerance parameter for model misspecification.
Figure 3: Comparison of ProBALL-UCB initialized with SOLD against LinUCB, mUCB, and mmUCB, for different choices of $\tau$ and confidence bound constructions. All variants of ProBALL-UCB perform no worse than LinUCB, and outperform mUCB and mmUCB. Shaded area depicts $1$-standard error confidence intervals over $30$ trials with fresh $\bm{\theta}$. The confidence intervals on regret thus account for the variation in frequentist regret for changing $\bm{\theta}$.
@@
Figure 3 depicts the result of this experiment. Once again, ProBALL-UCB performs no worse than LinUCB, no matter what we choose for $\tau$ and $\Delta_{\operatorname{off}}$, and the benefit of using tighter confidence bounds remains. With $\tau=0.5$, ProBALL-UCB with martingale Bernstein confidence bands stops using the projected estimates at around timestep $70$, but still continues to outperform LinUCB. Although mUCB and mmUCB perform slightly better than ProBALL-UCB and LinUCB at the beginning, the model misspecification incurred by discretizing the features into $d_{K}$ clusters ensures that it typically suffers linear regret in this scenario. The lower initial performance of ProBALL-UCB and Lin-UCB is a consequence of their higher initial exploration.

### v2

@@ 4 Offline Data Sharpens Online Optimism
1:Input: Projection matrix $\hat{\textbf{U}}\hat{\textbf{U}}^{\top}$, confidence bound $\Delta_{\operatorname{off}}$ from an offline uncertainty-aware method, e.g. SOLD.
@@
[…]
$\hat{\bm{\beta}}_{1,t}\leftarrow\textbf{U}(\textbf{U}^{T}\textbf{V}_{t}\textbf{U})^{-1}\textbf{U}^{\top}\textbf{b}_{t},\|\textbf{U}^{\top}(\bm{\beta}-\hat{\bm{\beta}}_{1,t})\|_{(\textbf{U}^{\top}\textbf{V}_{t}\textbf{U})^{-1}}\leq\alpha_{1,t}$
@@
$\|\bm{\beta}\|_{2}\leq R,\textbf{U}^{\top}\textbf{U}=\textbf{I}_{d_{K}},\textbf{U}\textbf{U}^{\top}\bm{\beta}=\bm{\beta},$
$\|\hat{\textbf{U}}^{\top}\textbf{U}\|_{F}\geq\sqrt{d_{K}-\Delta_{\operatorname{off}}^{2}/2}$
@@
We provide the following guarantee for LOCAL-UCB. Notice that our guarantee shows that for enough offline data with $N\gg T$, the effective dimension of the problem is $d_{K}$. It increases to $d_{A}$ as $T$ gets closer to $N$. The quality of the offline data is reflected in the coverage constants $\lambda_{\theta}$ and $\lambda_{A}$.
@@
Under Assumptions 1 and 2, if $\alpha_{1,t}=R\sqrt{\mu}+CR\sqrt{d_{K}\log(2T/\delta)}$ and $\alpha_{2,t}=R\sqrt{\mu}+CR\sqrt{d_{A}\log(2T/\delta)}$ for a universal constant $C$, then with probability at least $1-\delta$ over offline data and online rewards, LOCAL-UCB has regret $\operatorname{Reg}_{T}$ bounded by
$O\textstyle{\left(\min\left(Rd_{A}\sqrt{T},Rd_{K}\sqrt{T}\left(1+\frac{1}{\lambda_{\theta}\lambda_{A}^{3}}\sqrt{\frac{Td_{A}}{d_{K}N}}\right)\right)\right)}.$
@@
However, the subspace constraint $\|\hat{\textbf{U}}^{\top}\textbf{U}\|_{F}\geq\sqrt{d_{K}-\Delta_{\operatorname{off}}^{2}/2}$ is nonconvex. In fact, we lower bound a convex function in the constraint, making us search for $\widetilde{\textbf{U}}_{t}$ over a complicated star-shaped set. So, it is unclear if LOCAL-UCB can be made computationally efficient.
5 Lower Bound
We now establish that LOCAL-UCB is in fact minimax optimal. While we provide a full statement and proof of our lower bound in Appendix F, we provide an informal version here. Much like how we generate families of reward parameters in lower bound proofs for purely online regret, we are now generating a family of tuples of latent bandits (for the offline data) and reward parameters represented in the latent bandit (for the online interaction).
Theorem 3.
There exists a family of tuples $(F,\bm{\beta})$, where $F$ is a latent bandit with a rank $d_{K}$ latent subspace and $\bm{\beta}$ is a reward parameter in its support, so that for any offline behavior policy $\pi_{b}$ and any learner, $(i)$ $\lambda_{\theta}$ is uniformly bounded for all $F$, $(ii)$ there exists a $(F,\bm{\beta})$ such that the regret $\operatorname{Reg}(T,\bm{\beta})$ of the learner under offline data from $\pi_{b}$ and $F$ and online reward parameter $\bm{\beta}$ is bounded below by
$\Omega\left(\min\left(d_{A}\sqrt{T},d_{K}\sqrt{T}\left(1+\sqrt{\frac{d_{A}T}{d_{K}N}}\right)\right)\right)$
6 Practical Optimism with ProBALL-UCB
@@
Figure 1: Left: Geometric interpretations of LOCAL-UCB. Showing $\mathcal{C}^{t}_{\operatorname{on}}(\bm{\beta})\cap\mathcal{C}^{t}_{\operatorname{off}}(\bm{\beta})$ in green for three timepoints $t=t_{1},t_{2},t_{3}$. The dotted lines delineate the subspace confidence set. Right: Geometric interpretation of ProBALL-UCB. $\mathcal{C}^{t_{1}}_{\operatorname{on}}(\bm{\beta})\not\subset\widetilde{\mathcal{C}}^{t_{1}}_{\operatorname{off}}(\bm{\beta})$, so we continue to use projections; but by time $t_{2},$ $\mathcal{C}_{\operatorname{on}}^{t_{2}}(\bm{\beta})\subset\widetilde{\mathcal{C}}_{\operatorname{off}}^{t_{2}}(\bm{\beta})$, so we stop using projections.
While LOCAL-UCB is minimax-optimal, it is not computationally efficient due to the non-convex constraint discussed in Section 4. We address this by introducing ProBALL-UCB (Algorithm 3), a practical and computationally efficient algorithm. In this section, we first sketch the algorithm and then describe how it can be geometrically motivated as a relaxation of LOCAL-UCB.
ProBALL-UCB works in the subspace estimated by SOLD until the online confidence set is small enough. The algorithm maintains a low-dimensional confidence set, a high-dimensional confidence set, and swaps between them to achieve acceleration. Once the cumulative error of using the low-dimensional confidence set ($\approx\Delta_{\operatorname{off}}T$ in ProBALL-UCB) exceeds the cumulative error of using the high-dimensional confidence set ($\approx d_{A}\sqrt{T}$), we stop using the former. This is instantiated with LinUCB, but the same idea can be immediately applied to other algorithms like SupLinUCB or Bayesian algorithms like Thompson sampling.
We geometrically motivate our update rule here as a relaxation of LOCAL-UCB, and illustrate it in Figure 1, like in Section 4. We go through three stages of simplification over LOCAL-UCB, which suprisingly only leads to a minor degradation in provable guarantees.
  • Cruder offline confidence sets are used. We take the subspace estimated by SOLD, compute a point estimate for $\beta_{\star}$ within the subspace, and construct a ball that contains the LOCAL-UCB offline confidence set. The online confidence set is still the standard $d_{A}$-dimensional ellipsoid.
  • We wait for the offline confidence ball to contain the online confidence set, instead of taking intersections.
  • We use a computable proxy for this subset condition instead of explicitly checking it.
@@
$\operatorname{Reg}_{T}$ $=\textstyle{\widetilde{O}\left(\min\left(\operatorname{Reg}_{on,T},\operatorname{Reg}_{hyb,T}\right)\right)}.$
where $\operatorname{Reg}_{on,T}=Rd_{A}\sqrt{T}$ and $\operatorname{Reg}_{hyb,T}=Rd_{K}\sqrt{T}\left(1+\frac{1}{\lambda_{A}^{3}\lambda_{\theta}}\left(\sqrt{\frac{d_{A}T}{d_{K}N}}+\sqrt{\frac{d_{A}}{SN}\sum_{t=1}^{S}\kappa_{t}^{2}}\right)\right)$ In the worst case, $\kappa_{t}=O(t)$ and so $\frac{1}{S}\sum_{t=1}^{S}\kappa_{t}^{2}=O(T^{2})$, but if all features $\phi(x_{t},a_{t})$ lie in the span of $\hat{\textbf{U}}$ for $t\leq S$, then $\frac{1}{S}\sum_{t=1}^{S}\kappa_{t}^{2}=O(T)$.
@@ 7 Experiments
[…]
We now establish the practical efficacy of SOLD (Algorithm 1) and ProBALL-UCB (Algorithm 3) for linear latent contextual bandits through a series of numerical experiments. We perform a simulation study and a demonstration using real-life data. While specific details of the experiments and many ablation studies are in Appendix H, we sketch our experiments and discuss key observations in this section.
@@
In all experiments, we obtain confidence bounds $\Delta_{\operatorname{off}}$ using three different concentration inequalities – (1) Hoeffding as in Proposition 2 (H-ProBALL), (2) empirical Bernstein as in Proposition 1 (E-ProBALL), and (3) the martingale Bernstein concentration inequalities of Waudby-Smith and Ramdas, (2023) (M-ProBALL). We use a simpler expression for $\Delta_{\operatorname{off}}$, set $\tau^{\prime}=0$, and choose a suitable value of the hyperparameter $\tau$ to adjust for overly conservative $\Delta_{\operatorname{off}}$ [^5] Namely, we set $\Delta_{D}=0$ in $\Delta_{\operatorname{off}}$. Also, Lemma 6 and  some thought reveal that choosing $\tau^{\prime}=0$ recovers the ”good” version of ProBALL-UCB guarantees, if features are isotropic enough.. We later vary $\tau$ in ablation experiments to demonstrate that our results are not a consequence of our choice of hyperparameters. Finally, for the MovieLens experiments, we additionally design a natural Thompson sampling version of ProBALL-UCB to highlight the applicability of the ProBALL idea called ProBALL-TS in Appendix G. All experiments for ProBALL-TS are in Appendix H.3.2.
@@
Simulation study. We first perform a simulation study on a latent linear bandit with $d_{A}=50$ and $d_{K}=2$, with $5000$ trajectories generated offline. Further details are provided in Appendix H, and the results are presented in Figure 2. Note that ProBALL-UCB (Algorithm 3) performs no worse than LinUCB, no matter what we choose for $\tau$ and $\Delta_{\operatorname{off}}$. However, we see a clear benefit from using tighter confidence bounds – as $\Delta_{\operatorname{off}}$ gets smaller, Algorithm 3 chooses to utilize the projected estimate $\hat{\bm{\beta}}_{1,t}$ more often, resulting in better performance. Note that the kinks in the regret curves correspond to points where Algorithm 3 switches over to the higher dimensional optimism in step 7.
MovieLens dataset. In line with Hong et al., (2020), we assess the performance of our algorithms on real data using the MovieLens dataset. Like them, we filter the dataset to include only movies rated by at least $200$ users and vice versa, and apply probabilistic matrix factorization (PMF) to the rating matrix to generate ground truth user preferences for online experiments. Applying PMF gives $d_{K}=18$ and we choose $d_{A}=200$, generating 5000 trajectories offline. For baselines, we reproduce the methods of Hong et al., (2020, 2022) and implement LinUCB with canonical hyperparameter choices.
@@
We initialize ProBALL-UCB with a subspace estimated with an unregularized variant of SOLD (see Appendix G) that uses pseudo-inverses instead of inverses. This is due to difficulties in finding an appropriate regularization parameter for this large, noisy, and high-dimensional dataset. Figure 2 depicts the result of this experiment. Once again, ProBALL-UCB performs no worse than LinUCB, no matter what we choose for $\tau$ and $\Delta_{\operatorname{off}}$, and the benefit of using tighter confidence bounds remains. With $\tau=0.5$, ProBALL-UCB with martingale Bernstein confidence bands stops using the projected estimates at around timestep $70$, but still continues to outperform LinUCB. Although mUCB and mmUCB perform slightly better than ProBALL-UCB and LinUCB at the beginning, the model misspecification incurred by discretizing the features into $d_{K}$ clusters ensures that it typically suffers linear regret in this scenario. The lower initial performance of ProBALL-UCB and Lin-UCB is a consequence of their higher initial exploration.
Ablation study. While the end-to-end performance of ProBALL-UCB significantly improves over existing algorithms, we also address questions about various components of our method in Appendix H. First, we study the effect of varying the hyperparameter $\tau$ and note that our method stably outperforms existing methods at all reasonable values of $\tau$. Second, we compare different combinations of algorithms in our figures above, side by side. Finally, we evaluate the effect of offline data by plotting the online regret against the number of offline samples.
8 How General Are Latent Bandits?
While have established that latent bandits are a powerful framework for accounting for uncertainty in reward models, the extent of their generality is unclear. Are there other stateless decision processes that generalize over latent bandits? We cap off our contributions by establishing the generality of latent bandits. In this section, we show a de Finetti theorem for decision processes, demonstrating that every "coherent" and "exchangeable" stateless (contextual) decision process is a latent (contextual) bandit. We first define a stateless decision process at a high level of generality. [^6] Liu et al., (2023) work with a much more restrictive notion of a generalized bandit and use the original de Finetti theorem in some of their lemmas. See Appendix B.1 for a discussion.
Definition 2.
A stateless decision process (SDP) with action set $\mathcal{A}$ is a probability space $(\Omega,\mathcal{G},\mathbb{P})$ with a family of random maps $\mathcal{F}_{H}:\Omega\to(\mathcal{A}^{H}\to\mathbb{R}^{H})$ for $H\in\mathbb{N}\cup\{\infty\}$.
That is, given a sequence of actions $(a_{1},\dots a_{H})$, an SDP generates a random sequence of rewards $(Y_{1},\dots Y_{H})$. As such, we abuse notation to denote by $\mathcal{F}_{H}(a_{1},\dots a_{H})$ the random variable $\omega\mapsto\mathcal{F}_{H}(\omega)(a_{1},\dots a_{H})$. Without any coherence between $\mathcal{F}_{H}$ across $H$, a stateless process can behave arbitrarily for different horizons $H$. We present a natural coherence condition below, essentially requiring that a given action sequence should produce consistent rewards.
Definition 3.
A stateless decision process is coherent if for any $h\leq k\leq H,H^{\prime}\in\mathbb{N}\cup\{\infty\}$ and for any two action sequences $\tau,\tau^{\prime}$ of lengths $H$ and $H^{\prime}$ sharing the same actions $(a_{h},\dots a_{k})$ from index $h$ to $k$, with $\mathcal{F}_{H}(\tau)=(Y_{1},\dots Y_{H})$ and $\mathcal{F}_{H^{\prime}}(\tau^{\prime})=(Y_{1}^{\prime},\dots Y_{H^{\prime}}^{\prime})$, we have $(Y_{h},\dots Y_{k})=(Y_{h}^{\prime},\dots Y_{k}^{\prime})$, viewed as functions of $\Omega$.
It is natural to require equality in value and not just in distribution, since after taking an extra action, the values of past rewards stay the same, not just their distribution. For example, if we pull 10 different jackpot levers and then pull a new one, the previous 10 outcomes stay the same in value, not just in distribution. We also give a natural definition for exchangeability of a stateless decision process – namely that exchanging any two rewards should lead to the distribution obtained by exchanging the corresponding actions.
Definition 4.
A stateless decision process is exchangeable if for any permutation $\pi:[H]\to[H]$ and $\mathcal{F}_{H}(a_{1},\dots a_{H})=(Y_{1},\dots Y_{H})$, we have $\mathcal{F}_{h}(a_{\pi(1)},\dots a_{\pi(H)})\sim(Y_{\pi(1)},\dots Y_{\pi(H)})$.
Definition 5.
A latent bandit is a stateless decision process equipped with a random measure-valued function $F:\Omega\to(\mathcal{A}\to\mathcal{P}(\mathbb{R}))$ so that for any $H$ and action sequence $(a_{1},\dots a_{H})$, the rewards $(Y_{1},\dots Y_{H}):=\mathcal{F}_{H}(a_{1},\dots a_{H})$ are independent conditioned on $F$. Moreover, the conditional distribution $\mathcal{L}[Y_{h}\mid F]=F(a_{h})$ for all $h\leq H$. [^7] We abuse notation twice here. First, we write $F(a_{h}):=(\omega\mapsto F(\omega)(a_{h}))$. Second, as the regular conditional distribution $\mathcal{L}[Y_{h}\mid F]$ is a kernel that maps from $\Omega\times\mathcal{B}\to\mathbb{R}$, we view $F(a_{h})$ as its curried map $(\omega,B)\mapsto F(a_{h})(\omega)(B)$. A discussion of issues (measurability and well-definedness) is in Appendix B.2.
While exchangeability and coherence are reasonable conditions on an SDP and are clearly satisfied by latent bandits, it is a-priori unclear if they are sufficient to ensure that the SDP is a latent bandit. As only exchanging rewards from the same action preserves the distribution, standard de Finetti proof ideas do not immediately apply. After all, it is possible that an SDP could be cleverly designed to choose rewards adaptively across time and satisfy these properties. Reassuringly, no such counterexamples exist, guaranteed by the following theorem.
Theorem 5 (De Finetti Theorem for Stateless Decision Processes).
We show in Lemma 2 in Appendix A that coherence is not a consequence of exchangeability – it is a necessary condition for being a latent bandit. Finally, we analogously consider contexts and define "transition-agnostic contextual decision processes" (TACDPs) in Appendix B.2. We define coherence and coherence and exchangeability for TACDPs, and define latent contextual bandits by simply replacing $\mathcal{A}$ with $\mathcal{X}\times\mathcal{A}$ in the definitions above. We then show an analogous de Finetti theorem, as a corollary of our proof of Theorem 5. See Appendix B.6 for more details.
Finally, note that this section is faithful to the rest of this paper. A linear latent contextual bandit is a latent contextual bandit where the random measure-valued function $F:\Omega\to((\mathcal{X}\times\mathcal{A})\to\mathcal{P}(\mathbb{R}))$ is defined by setting $F(\omega)(x,a)$ to be the distribution given by $\phi(x,a)^{\top}\textbf{U}_{\star}\bm{\theta}(\omega)+\epsilon$, for any $\omega\in\Omega$ and $(x,a)\in(\mathcal{X}\times\mathcal{A})$.